# Scale dependence improvement of the quartic scalar field thermal effective potential in the optimized perturbation theory

### v1

@@ Appendix D VRG Effective Potential
- $\left.J_{1,\bar{\Omega}}\right)\left.\left.+\left(\frac{C_{1}}{3}+C_{2}+\frac{L_{\bar{\mu}}}{6}\right)\bar{\Omega}^{2}\right]\right\}+\frac{\lambda^{2}\varphi^{4}}{48\xi^{2}(4\pi)^{2}}(2-3J_{2,\bar{\Omega}})+F_{\rm vac,2}+F_{\varphi},$
where
$F_{\varphi}$ $=$ $\lambda^{2}\varphi^{2}\left\{\frac{m^{2}(19(\xi-1)-34\ln\xi)}{216\xi^{7/3}(4\pi)^{4}}\left(L_{\bar{\mu}}+J_{2,\bar{\Omega}}\right)+\frac{m^{2}(19(\xi-1)-34\ln\xi)}{432\xi^{7/3}(4\pi)^{4}}\left(1-\frac{\eta^{2}}{T\bar{\Omega}}J_{2,\bar{\Omega}}^{\prime}\right)\right$.
$+$ $\frac{\eta^{2}(\xi+34\ln\xi-1)\left(L_{\bar{\mu}}+J_{2,\bar{\Omega}}\right)}{72\xi^{2}(4\pi)^{4}}-\frac{(\xi+34\ln\xi-1)\left[(L_{\bar{\mu}}+1)\bar{\Omega}^{2}-T^{2}J_{1,\bar{\Omega}}\right]}{72\xi^{2}(4\pi)^{4}}$
- $\frac{m^{2}(2\xi-17\ln\xi-2)(L_{\bar{\mu}}+1)}{27\xi^{7/3}(4\pi)^{4}}+\frac{17\eta^{2}\ln\xi}{36\xi^{2}(4\pi)^{4}}+\frac{(\xi-1)\bar{\Omega}^{2}}{72\xi^{2}(4\pi)^{4}}+\frac{\lambda^{2}\varphi^{4}\left(\xi-1+17\ln\xi\right)}{216\xi^{2}(4\pi)^{2}}$
$+$ $\frac{m^{2}}{\xi^{7/3}(4\pi)^{4}}\left[-\frac{17(32\xi+103)\ln\xi}{5832}+\frac{289\ln^{2}\xi}{729}\right.+(\xi(3888\zeta(3)+973)-3888\zeta(3)-\left.\left.3997)\frac{(\xi-1)}{11664}\right]\right\},$
and
$F_{\rm vac,2}$ $=$ $-\frac{m^{2}\eta^{4}\lambda\left(19(1-\xi)-34\ln\xi\right)}{216(4\pi)^{4}\,\xi^{4/3}\bar{\Omega}^{2}}\left[1-\frac{\bar{\Omega}}{2T}J_{2,\bar{\Omega}}^{\prime}\right]+\frac{\lambda^{2}m^{4}}{\xi^{8/3}(4\pi)^{6}}\left[-\frac{1}{60}\pi^{4}\xi^{8/3}+\frac{6337\xi^{8/3}}{1440}+\frac{\pi^{4}\xi^{3}}{90}\right$. (74)
- $\frac{3351853\xi^{3}}{1889568}+\frac{\pi^{4}\xi^{2}}{180}-\frac{3522535\xi^{2}}{1259712}+\frac{5\zeta(5)}{18}+\frac{42653\xi^{2}\ln(\xi)}{314928}-\frac{109649\xi}{196830}+\frac{24565\ln^{3}(\xi)}{118098}$
- $\frac{5491\xi\ln^{2}(\xi)}{39366}-\frac{2023\ln^{2}(\xi)}{78732}-\frac{106301\xi\ln(\xi)}{157464}+\frac{140743\ln(\xi)}{157464}+\frac{229}{40}\xi^{8/3}\zeta(3)+\frac{15}{2}\xi^{8/3}\zeta(5)$
- $\frac{577\xi^{3}\zeta(3)}{162}-\frac{40\xi^{3}\zeta(5)}{9}-\frac{113\xi^{2}\zeta(3)}{54}-\frac{10\xi^{2}\zeta(5)}{3}+\frac{34}{81}\xi^{2}\zeta(3)\ln(\xi)-\frac{76\xi\zeta(3)}{135}$
- $\frac{68}{81}\xi\zeta(3)\ln(\xi)+\frac{319\zeta(3)}{648}+\frac{2745749}{3779136}+\left.\frac{34}{81}\zeta(3)\ln(\xi)\right]-\frac{17\lambda^{2}\ln(\xi)}{72\xi^{2}(4\pi)^{6}}\left[\eta^{4}-T^{4}J_{1,\bar{\Omega}}^{2}\right$.
- $2\eta^{2}T^{2}J_{1,\bar{\Omega}}J_{2,\bar{\Omega}}-2\eta^{2}\left(L_{\bar{\mu}}^{2}+L_{\bar{\mu}}+1\right)\bar{\Omega}^{2}+\left.\left(L_{\bar{\mu}}^{2}+2L_{\bar{\mu}}+2\right)\bar{\Omega}^{4}\right]+\frac{\lambda^{2}m^{2}\kappa_{1}}{432\xi^{7/3}(4\pi)^{6}}\left[\right$.
$+$ $2\eta^{2}J_{2,\bar{\Omega}}^{2}+TJ_{1,\bar{\Omega}}\left(2TJ_{2,\bar{\Omega}}-\frac{\eta^{2}J_{2,\bar{\Omega}}^{\prime}}{\bar{\Omega}}\right)+\left.\eta^{2}\left(-2L_{\bar{\mu}}^{2}+2L_{\bar{\mu}}+1\right)+\left(2L_{\bar{\mu}}^{2}+2L_{\bar{\mu}}+1\right)\bar{\Omega}^{2}\right]$
- $\frac{\lambda^{2}m^{2}\kappa_{2}}{23328\xi^{7/3}(4\pi)^{6}}\left[(L_{\bar{\mu}}+1)\bar{\Omega}^{2}-T^{2}J_{1,\bar{\Omega}}\right.-\left.\frac{\eta^{4}}{2\bar{\Omega}^{2}}\left(1-\frac{\bar{\Omega}J_{2,\bar{\Omega}}^{\prime}}{2T}\right)-\eta^{2}\left(L+J_{2,\bar{\Omega}}\right)\right]$
$+$ $\frac{\lambda^{2}m^{4}\kappa_{1}^{2}}{93312\xi^{8/3}(4\pi)^{6}}\left[-\frac{4\eta^{2}\left(\eta^{2}+2\bar{\Omega}^{2}\right)}{\bar{\Omega}^{4}}-\frac{\eta^{4}J_{2,\bar{\Omega}}^{\prime\prime}}{T^{2}\Omega^{2}}\right.+\left.\frac{\eta^{2}\left(\eta^{2}+4\bar{\Omega}^{2}\right)J_{2,\bar{\Omega}}^{\prime}}{T\bar{\Omega}^{3}}-8\left(L+J_{2,\bar{\Omega}}\right)\right],$
where we have also defined the quantities $\kappa_{1}$ and $\kappa_{2}$ as
$\kappa_{1}=19(\xi-1)-34\ln\xi,$ (75)
and
$\kappa_{2}$ $=$ $3(\xi-1)(2584\ln\xi+[2592\zeta(3)-367]\xi$ (76)
- $2592\zeta(3)-3269)+180\kappa_{1}+8\kappa_{1}^{2}.$

### v2

@@ Appendix E VRG Effective Potential
$+$ $\left.(\xi-1)(171L_{\bar{\mu}}-2\xi(648\zeta(3)+179)+1296\zeta(3)+1435)\right]+\bar{\Omega}^{2}\left[831060L_{\bar{\mu}}^{2}(\xi-1)-981264\xi+5923585\right$.
- $34680(243L_{\bar{\mu}}+76\xi+257)\ln^{2}\xi-314928\pi^{4}\xi^{8/3}+83154114\xi^{8/3}+209952\pi^{4}\xi^{3}-33518530\xi^{3}$
$+$ $104976\pi^{4}\xi^{2}-54577905\xi^{2}+3930400\ln^{3}\xi-9720L_{\bar{\mu}}(\xi-1)(\xi(648\zeta(3)+179)-648\zeta(3)-803)$
$+$ $108177768\xi^{8/3}\zeta(3)+141717600\xi^{8/3}\zeta(5)-67301280\xi^{3}\zeta(3)-83980800\xi^{3}\zeta(5)-45839520\xi^{2}\zeta(3)$
- $62985600\xi^{2}\zeta(5)+1959552\xi\zeta(3)++3003480\zeta(3)+5248800\zeta(5)-1020\ln\xi\left(10206L_{\bar{\mu}}^{2}-162L_{\bar{\mu}}(19\xi-97)\right$.
- $\left.\xi^{2}(7776\zeta(3)+2509)+4\xi(3888\zeta(3)+2357)-7776\zeta(3)-5218\right),$